# Xing Huang(a), Eva Kopfer(b), Panpan Ren(c) a)

### v3

@@ Abstract
The exponential ergodicity in the $L^{1}$-Wasserstein distance for partially dissipative McKean-Vlasov SDEs has been extensively studied. However, the question of exponential ergodicity in the $L^{2}$-Wasserstein distance and relative entropy has remained unresolved. This paper addresses the problem by establishing the log-Sobolev inequality for both the time-marginal distributions and the invariant probability measure, providing a positive resolution. As part of the groundwork, the log-Sobolev inequality is investigated for the associated time-inhomogeneous semigroup. The main results are further extended to degenerate diffusion.
@@ 1 Introduction
We aim to advance the study of functional inequalities in McKean–Vlasov SDEs and their corresponding decoupled equations. The McKean–Vlasov SDE, first introduced by McKean [23], was developed to study the nonlinear Fokker-Planck equation. The idea goes back to Kac [18, 19], who characterized nonlinear PDEs in Vlasov’s kinetic theory. The main tool is the so-called “propagation of chaos” which builds a corner stone in mean field interacting particle systems, see e.g. [9, 24, 31]. The McKean–Vlasov SDE is the prototype of a distribution–dependent stochastic differential equation (DDSDE).
In general, the McKean–Vlasov SDE is given by
@@
where $\{W_{t}\}_{t\geq 0}$ is an $n$-dimensional Brownian motion on some complete filtered probability space $(\Omega,\mathscr{F},(\mathscr{F}_{t})_{t\geq 0},\mathbb{P})$ and $\mathscr{L}_{X_{t}}$ is the distribution of $X_{t}$. The coefficients $b:[0,\infty)\times\mathbb{R}^{d}\times\mathscr{P}(\mathbb{R}^{d})\rightarrow\mathbb{R}^{d}$, $\sigma:[0,\infty)\times\mathbb{R}^{d}\times\mathscr{P}(\mathbb{R}^{d})\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ are assumed to be measurable and bounded on bounded sets, where $\mathscr{P}(\mathbb{R}^{d})$ is the set of all probability measures on $\mathbb{R}^{d}$ equipped with the weak topology.
In general, we will consider the solution of (1.1) in a subspace of $\mathscr{P}(\mathbb{R}^{d})$. To this end, for $p\geq 1$, let
@@
The well-posedness for (1.1) in the weak and strong sense has been intensively investigated, see e.g. [36, 39] and references within. When (1.1) is well-posed in $\mathscr{P}_{p}(\mathbb{R}^{d})$, for given $\gamma\in\mathscr{P}_{p}(\mathbb{R}^{d})$ we denote $P_{t}^{*}\gamma=\mathscr{L}_{X_{t}}$ with $\mathscr{L}_{X_{0}}=\gamma$.
@@
for some $\alpha>\beta>0$, exponential ergodicity holds in the $L^{2}$-Wasserstein distance as well as in relative entropy, see [17, 30, 36] and references therein. In the partially dissipative case, i.e. dissipation only occurs at long distances, exponential ergodicity has been established in the $L^{1}$-Wasserstein distance, see [11, 16, 21, 39] and references therein, using asymptotic reflection couplings. Additionally, in the symmetric case, i.e. the invariant probability measure has explicit representation, exponential ergodicity in the $L^{2}$-Wasserstein distance has been proven under the condition that the drift is bounded in the distribution variable and the disffusion is non-degenerate and only depends on the distribution variable, see [45]. For recent progresses on long-time behaviour of McKean-Vlasov SDEs, please refer to [43] on exponential ergodicity in total variation distance, [14] concerning exponential ergodicity in $L^{2}$-Wasserstein distance and mean field entropy for kinetic McKean-Vlasov SDEs, and [26] regarding non-asymptotic entropic bounds of stochastic algorithms.
Our main results show exponential ergodicity in both the $L^{2}$-Wasserstein distance and relative entropy for partially dissipative drifts, possibly unbounded in the distribution variable, with non-degenerate or degenerate diffusion. This further expands the results in [30] towards partially dissipative drifts and the results in [4, 32] towards drifts, which are not of gradient-type.
As a main tool in the non-degenerate case, we will use Wang’s observation in [41], which states that exponential ergodicity in relative entropy can be deduced from log-Sobolev inequalities. The study of log-Sobolev inequalities originated in the works of Leonard Gross [12], who introduced this for Gaussian measures. The seminal work by Bakry and Émery [1] links the log-Sobolev inequality to lower curvature bounds of diffusion processes via the so-called $\Gamma$-calculus introduced therein. Since then, log-Sobolev inequalities are an integral part in the study of SDEs and PDEs, e.g. in [6, 10, 13, 20, 22, 27, 33].
@@
(1.2) $\text{\rm{d}}X_{s,t}^{\mu,x}=b_{t}(X_{s,t}^{\mu,x},\mu_{t})\text{\rm{d}}t+\sigma_{t}(X_{s,t}^{\mu,x},\mu_{t})\text{\rm{d}}W_{t},\ \ X_{s,s}^{\mu,x}=x\in\mathbb{R}^{d},t\geq s\geq 0,$
@@
(1.3) $P_{s,t}^{\mu}f(x)=\mathbb{E}f(X_{s,t}^{\mu,x}),\ \ t\geq s\geq 0,f\in\mathscr{B}_{b}(\mathbb{R}^{d}).$
These results apply to both the non-degenerate and degenerate case.
In order to obtain the log-Sobolev inequality for the invariant probability measure, we focus on the time-homogeneous case and use the hyperboundedness derived from Wang’s Harnack inequality. This improves the results in [30] towards partially dissipative drifts and the results in [32, 4] towards drifts, which are not of gradient-type.
The remainder of this paper is organized as follows. In Section 2, we study the non-degenerate case and we prove the log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}$ (Theorem 2.1) under the monotonicity assumption (A). In the partially dissipative setting (H), we derive the log-Sobolev inequality for the invariant probability measure (Theorem 2.6). As an application, we deduce the exponential ergodicity in $\mathbb{W}_{2}$ and relative entropy in Theorem 2.7. In Section 3, we concentrate on the degenerate case. We establish the log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}$ under the Lipschitz condition (B). For partially dissipative drifts satisfying ((C)), we derive the log-Sobolev inequality for the invariant probability measure in Theorem 3.5. As an application, we obtain exponential ergodicity in $\mathbb{W}_{2}$ and relative entropy for both uniformly dissipative (Theorem 3.7) and partially dissipative drifts (Theorem 3.9).
2 Non-degenerate case
@@ 2.1 Log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}\mu_{0}$
We will adopt the $\Gamma$-calculus developed in [1] to our SDE setting. To this end, for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$, we define
@@
The associated diffusion operator is then given by
(2.1) $\mathscr{L}_{t}^{\mu}f=\frac{1}{2}\mathrm{tr}(\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast}\nabla^{2}f)+\nabla_{b_{t}^{\mu}}f,\ \ f\in C_{0}^{\infty}(\mathbb{R}^{d}),$
@@
We make the following assumptions:
  1. (A)
    For any $t\geq 0$, $b_{t}$ and $\sigma_{t}$ are continuous in $\mathbb{R}^{d}\times\mathscr{P}_{2}(\mathbb{R}^{d})$. There exist constants $K_{1},K_{I}>0$ such that for all $t\geq 0$, $x,y\in\mathbb{R}^{d}$ and $\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{d})$,
    (2.2) $\leq K_{1}|x-y|^{2}+K_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2}.$
Here, $K_{I}$ represents the dependence of $b$ and $\sigma$ on measure variable.
Under (A), (1.1) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{d})$ and recall that $P_{t}^{\ast}\gamma$ is the distribution of the solution to (1.1) with initial distribution $\gamma\in\mathscr{P}_{2}(\mathbb{R}^{d})$. Moreover, for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$, (1.2) has a unique solution $X_{s,t}^{\mu,x}$ with associated time-inhomogenous semigroup $P_{s,t}^{\mu}$. See e.g. [39, Theorem 3.3.1] for these facts.
@@
Assume $({\bf A})$ and suppose $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$ satisfying for some constant $\delta_{1}>0$,
(2.3) $\sigma_{t}(\gamma)\sigma_{t}(\gamma)^{\ast}\leq\delta_{1}I_{d\times d},\ \ \gamma\in\mathscr{P}_{2}(\mathbb{R}^{d}).$
Then, $P_{s,t}^{\mu}$ satisfies the log-Sobolev inequality: for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$ and $0<f\in C_{0}^{\infty}(\mathbb{R}^{d})$, the smooth functions with compact support on $\mathbb{R}^{d}$,
(2.4) $P_{s,t}^{\mu}(f\log f)-P_{s,t}^{\mu}(f)\log P_{s,t}^{\mu}(f)\leq\frac{2\delta_{1}}{K_{1}}(\text{\rm{e}}^{K_{1}(t-s)}-1)P_{s,t}^{\mu}\big{(}|\nabla f^{\frac{1}{2}}|^{2}\big{)},\ \ 0\leq s\leq t.$
If moreover $\mu_{0}$ satisfies the log-Sobolev inequality: for some $L_{0}\geq 0$,
[…]
then $P_{t}^{\ast}\mu_{0}$ satisfies the log-Sobolev inequality:
(2.6) $\begin{split}(P_{t}^{\ast}\mu_{0})&(f\log f)-(P_{t}^{\ast}\mu_{0})(f)\log(P_{t}^{\ast}\mu_{0})(f)\\ &\leq 2\Big{(}L_{0}\text{\rm{e}}^{K_{1}t}+\frac{\delta_{1}}{K_{1}}(\text{\rm{e}}^{K_{1}t}-1)\Big{)}(P_{t}^{\ast}\mu_{0})\big{(}|\nabla f^{\frac{1}{2}}|^{2}\big{)},\ \ 0<f\in C_{0}^{\infty}(\mathbb{R}^{d}),t\geq 0.\end{split}$
@@
(1) By an approximation trick, we may and do assume that $b_{t}(\cdot,\gamma)$ is smooth. Recall that $X_{s,t}^{\mu,x}$ and $P_{s,t}^{\mu}$ is defined in (1.2) and (1.3). By ((A)) and $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$, it is standard to derive that
$|X_{s,t}^{\mu,x}-X_{s,t}^{\mu,y}|\leq\text{\rm{e}}^{\frac{1}{2}K_{1}(t-s)}|x-y|,\ \ x,y\in\mathbb{R}^{d},0\leq s\leq t.$
This implies that for any $x\in\mathbb{R}^{d},0\leq s\leq t,f\in C_{0}^{\infty}(\mathbb{R}^{d})$,

### v4

@@ Abstract
The exponential ergodicity of partially dissipative McKean-Vlasov SDEs in the $L^{1}$-Wasserstein distance has been extensively studied using asymptotic reflection coupling. However, the reflection coupling method is not applicable for the exponential ergodicity in $L^{2}$-Wasserstein distance and relative entropy. In this paper, we first establish uniform log-Sobolev inequalities (in the frozen measure variable with bounded second moments) for the invariant probability measure of the corresponding SDEs with frozen distribution. Second, for the McKean-Vlasov SDEs, we combine the log-Harnack inequality and Talagrand’s inequality to derive exponential ergodicity in both $L^{2}$-Wasserstein distance and relative entropy. Furthermore, we extend these main results to the case of degenerate diffusion.
@@ 1 Introduction
1.1 Backgroud on McKean-Vlasov SDEs
The McKean-Vlasov SDE, first introduced by McKean [32], was developed to study the nonlinear Fokker-Planck equation. The idea goes back to Kac [26, 27], who characterized nonlinear PDEs in Vlasov’s kinetic theory. The main tool is the so-called “propagation of chaos” which builds a corner stone in mean field interacting particle systems, see e.g. [11, 33, 42]. In general, the McKean-Vlasov SDE is given by
@@ 1.1 Backgroud on McKean-Vlasov SDEs
where $\{W_{t}\}_{t\geq 0}$ is an $n$-dimensional Brownian motion on some complete filtered probability space $(\Omega,\mathscr{F},(\mathscr{F}_{t})_{t\geq 0},\mathbb{P})$ and $\mathscr{L}_{X_{t}}$ is the distribution of $X_{t}$. The coefficients $b:[0,\infty)\times\mathbb{R}^{d}\times\mathscr{P}(\mathbb{R}^{d})\rightarrow\mathbb{R}^{d}$, $\sigma:[0,\infty)\times\mathbb{R}^{d}\times\mathscr{P}(\mathbb{R}^{d})\rightarrow\mathbb{R}^{d}\otimes\mathbb{R}^{n}$ are assumed to be measurable and bounded on bounded sets, where $\mathscr{P}(\mathbb{R}^{d})$ is the set of all probability measures on $\mathbb{R}^{d}$ equipped with the weak topology. We will consider the solution of (1.1) in a subspace of $\mathscr{P}(\mathbb{R}^{d})$. To this end, for $p\geq 1$, let
@@
When (1.1) is well-posed in $\mathscr{P}_{p}(\mathbb{R}^{d})$, for given $\gamma\in\mathscr{P}_{p}(\mathbb{R}^{d})$ we denote $P_{t}^{*}\gamma=\mathscr{L}_{X_{t}}$ with $\mathscr{L}_{X_{0}}=\gamma$.
@@
In this paper, we aim to advance the study of functional inequalities for McKean-Vlasov SDEs and their associated equations with frozen distribution and to apply these results to analyze exponential ergodicity in both the $L^{2}$-Wasserstein distance and relative entropy. Our approach builds on suitable Wang’s Harnack inequality and Wang’s observation in [52](and [43] in the kinetic case), which states that exponential ergodicity in relative entropy can be deduced from log-Sobolev inequalities. To place our results into context, we briefly review key results on McKean-Vlasov SDEs, focusing on log-Sobolev inequalities and exponential ergodicity under various metrics.
1.2 Related works
1.2.1 Log-Sobolev inequalities for McKean-Vlasov SDEs
We begin by reviewing existing results on log-Sobolev inequalities. The study of log-Sobolev inequalities originated in the works of Leonard Gross [19], who introduced this for Gaussian measures. The seminal work by Bakry and Émery [1] links the log-Sobolev inequality to lower curvature bounds of diffusion processes via the so-called $\Gamma$-calculus introduced therein. Since then, log-Sobolev inequalities are an integral part in the study of SDEs and PDEs, e.g. in [6, 15, 21, 29, 31, 37, 44].
Recently, the log-Sobolev inequality for McKean-Vlasov SDEs also attracts much attention. For SDEs of the form
$\text{\rm{d}}X_{t}=b(X_{t},\mathscr{L}_{X_{t}})\,\text{\rm{d}}t+\sigma\,\text{\rm{d}}W_{t},$
[28, Corollary 2.7] establishes a uniform-in-time log-Sobolev inequality. They assume that $\sigma=I_{d\times d}$,
$b(x,\mu)=-\nabla U^{1}(x)+\int_{\mathbb{R}^{d}}-\nabla U^{2}(x-y)\,\mu(\text{\rm{d}}y),$
$U^{1},U^{2}$ are twice continuously differentiable, $\nabla^{2}(U^{1}+U^{2})\geq\alpha I_{d\times d}$ for some $\alpha>0$, $\nabla U^{2}$ is even and bounded or Lipschitz, and $\nabla U^{i}$ satisfies certain monotonicity and growth conditions. The result holds provided the Lipschitz constant of $\nabla U^{2}$ is smaller than $\alpha$ and the initial law $\mathscr{L}_{X_{0}}$ satisfies a log-Sobolev inequality.
On the torus $\mathbb{T}^{d}$, similar results are obtained under convolution-type interactions. In [28, Corollary 2.9], when $b(x,\mu)=\int_{\mathbb{T}^{d}}K(x-y)\,\mu(\text{\rm{d}}y)$ for a Lipschitz kernel $K$, a uniform log-Sobolev inequality holds if $\|\mathrm{div}\,K\|_{\infty}$ is sufficiently small and the initial density has bounded derivatives and is uniformly bounded away from zero and infinity. [20, Theorem 2.5] extends this by allowing divergence-free interaction kernels of the form $K=\mathrm{div}\,V$, again under smoothness and boundedness conditions on the initial density, using the Holley-Stroock perturbation lemma and the log-Sobolev inequality of the uniform distribution on $\mathbb{T}^{d}$.
For time-inhomogeneous SDEs, [35, Theorem 1.3] shows that a uniform-in-time log-Sobolev inequality holds when the drift $b_{t}$ satisfies a local monotonicity condition inside a ball, a strong dissipativity condition outside, and a mild growth condition, provided the diffusion coefficient $\sigma^{2}$ is large enough. This result is then applied to McKean-Vlasov SDEs in [35, Proposition 3.9] with drifts of the form $\nabla U-\int\nabla_{x}W(\cdot,y)\,\mathscr{L}_{X_{t}}(\text{\rm{d}}y)$, where $W$ is split into a bounded component and a Lipschitz component with sufficiently small Lipschitz constant. Under these assumptions, the uniform-in-time log-Sobolev inequality holds if it holds for the initial law.
In this paper, Theorem 2.1 improves upon existing results by establishing a explicit log-Sobolev inequality for time-inhomogeneous semigroup with frozen distribution and McKean-Vlasov SDEs under mild assumptions: the drift is only locally bounded and continuous in the spatial variable, and the diffusion may depend on the measure. This is different from prior works that require $C^{2}$ regularity on the coefficients, see for instance [8, Theorem 4.2]. When the coefficients satisfy partially dissipative condition, both in the non-degenerate and degenerate case, the invariant probability measure for the time-homogeneous SDEs with frozen distribution also satisfies log-Sobolev inequality even if the invariant probability measure does not have explicit representation. The key point also lies in that this log-Sobolev inequality is uniform in the froze measure with bounded second moment.
1.2.2 Exponential ergodicity for McKean-Vlasov SDEs
@@
for some $\alpha>\beta>0$, exponential ergodicity holds in the $L^{2}$-Wasserstein distance as well as in relative entropy, see [25, 41, 47] and references therein. In the partially dissipative case, i.e. dissipation only occurs at long distances, exponential ergodicity has been established in the $L^{1}$-Wasserstein distance, see [18, 24, 30, 50] and references therein, using asymptotic reflection couplings. Additionally, in the symmetric case, i.e. the invariant probability measure has explicit representation, exponential ergodicity in the $L^{2}$-Wasserstein distance has been proven under the condition that the drift is bounded in the distribution variable and the diffusion is non-degenerate and only depends on the distribution variable, see [56]. For recent progresses on long-time behaviour of McKean-Vlasov SDEs, we refer to [54] on exponential ergodicity in total variation distance, [22] concerning exponential ergodicity in $L^{2}$-Wasserstein distance and mean field entropy for kinetic McKean-Vlasov SDEs under non-convex assumption and [36] regarding non-asymptotic entropic bounds of stochastic algorithms. Also relevant is [34, Theorem 1.2], which proves $L^{2}$-Wasserstein contraction for distribution-independent SDEs. For this result, the drift is partially dissipative, and $\sigma=\sigma_{0}I_{d\times d}$ for large enough $\sigma_{0}$. The author’s method relies on constructing a new distance that is equivalent to the squared Euclidean norm and then employing a synchronous coupling argument.
This paper establishes exponential ergodicity in the $L^{2}$-Wasserstein distance and relative entropy for McKean-Vlasov SDEs with partially dissipative drifts and weak enough interaction, which may be unbounded in the distribution variable and encompass both non-degenerate and degenerate diffusions. This finding extends the scope of [41] to partially dissipative settings and generalizes [4, 43] to non-gradient-type drifts. The proof relies on uniform log-Sobolev inequalities for the frozen system, derived using Wang’s Harnack inequality, the semigroup log-Sobolev inequality, and the weak Poincaré inequality. Most importantly, Wang’s Harnack inequality allows us to skip the explicit form of the invariant probability measure.
Outline
The remainder of this paper is organized as follows. In Section 2, we study the non-degenerate case and we prove the log-Sobolev inequality for semigroup $P_{s,t}^{\mu}$ defined in (2.2) and $P_{t}^{\ast}$ in Theorem 2.1 under the monotonicity assumption (A). In the partially dissipative setting (H), we derive the log-Sobolev inequality for the invariant probability measure associated to the time-homogeneous SDEs with frozen distribution (Theorem 2.11). As an application, we deduce the exponential ergodicity in $L^{2}$-Wasserstein distance and relative entropy for McKean-Vlasov SDEs in Theorem 2.13. In Section 3, we concentrate on the degenerate case. For partially dissipative drifts, we derive the log-Sobolev inequality for the invariant probability measure associated to the time-homogeneous SDEs with frozen distribution in Theorem 3.5. As an application, we obtain exponential ergodicity for McKean-Vlasov SDEs in $L^{2}$-Wasserstein distance and relative entropy for both uniformly dissipative (Theorem 3.7) and partially dissipative drifts (Theorem 3.9). In Section 4, we provide the well-posedness of McKean-Vlasov SDEs and SDEs with frozen distribution under the condition (A) in Theorem 4.1. The approximation approaches used to prove Theorem 2.1 also have been provided.
2 Non-degenerate case
@@
(2.1) $\text{\rm{d}}X_{s,t}^{\mu}=b_{t}(X_{s,t}^{\mu},\mu_{t})\text{\rm{d}}t+\sigma_{t}(X_{s,t}^{\mu},\mu_{t})\text{\rm{d}}W_{t},\ \ t\geq s\geq 0,$
@@
(2.2) $P_{s,t}^{\mu}f(x)=\mathbb{E}f(X_{s,t}^{\mu,x}),\ \ t\geq s\geq 0,f\in\mathscr{B}_{b}(\mathbb{R}^{d}),$
where $X_{s,t}^{\mu,x}$ solve (2.1) with $X_{s,s}^{\mu,x}=x\in\mathbb{R}^{d}$.
  1. (A)
    $b:[0,\infty)\times\mathbb{R}^{d}\times\mathscr{P}_{2}(\mathbb{R}^{d})\mapsto\mathbb{R}^{d}$ is bounded on bounded sets. For any $t\geq 0$ and $\gamma\in\mathscr{P}_{2}(\mathbb{R}^{d})$, $b_{t}(\cdot,\gamma)$ and $\sigma_{t}(\cdot,\gamma)$ are continuous in $\mathbb{R}^{d}$. There exist constants $K_{1},K_{I}>0$ such that for all $t\geq 0$, $x,y\in\mathbb{R}^{d}$ and $\gamma,\tilde{\gamma}\in\mathscr{P}_{2}(\mathbb{R}^{d})$,
    (2.3) $\leq K_{1}|x-y|^{2}+K_{I}\mathbb{W}_{2}(\gamma,\tilde{\gamma})^{2},$
    and
    $\|\sigma_{t}(x,\gamma)\|_{HS}^{2}\leq K_{1}(1+|x|^{2}+\gamma(|\cdot|^{2})).$
Note that in (A), the local Lipschitz continuity of $b$ and $\sigma$ in spatial variable is not required. Under (A), (1.1) is well-posed in $\mathscr{P}_{2}(\mathbb{R}^{d})$ and recall that $P_{t}^{\ast}\gamma$ is the distribution of the solution to (1.1) with initial distribution $\gamma\in\mathscr{P}_{2}(\mathbb{R}^{d})$. Moreover, for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$, (2.1) has a unique solution $X_{s,t}^{\mu,x}$ with initial value $x\in\mathbb{R}^{d}$ and associated time-inhomogenous semigroup $P_{s,t}^{\mu}$. More details please refer to Theorem 4.1 in appendix.
@@ 2.1 Log-Sobolev inequality for $P_{s,t}^{\mu}$ and $P_{t}^{\ast}\mu_{0}$
In this subsection, we will establish a log-Sobolev inequality for the semigroup $P_{s,t}^{\mu}$ defined in (2.2). When $b,\sigma$ are distribution free, a related result was previously obtained in [8, Theorem 4.2] under condition ((A)) and suitable $C^{2}$-regularity on the coefficients, using classical stochastic methods. Here, we derive this inequality under slightly different assumptions by employing the $\Gamma$-calculus developed in [1], together with an approximation argument.
To this end, for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$, we define
@@
Let $C_{b}^{\infty}(\mathbb{R}^{d}):=\{f\in C^{\infty}(\mathbb{R}^{d}):\|\nabla^{i}f\|_{\infty}<\infty,i\geq 1\}$, and $C_{0}^{\infty}(\mathbb{R}^{d})$ denote the set of all smooth functions on $\mathbb{R}^{d}$ with compact support. The associated diffusion operator is then given by
(2.4) $\mathscr{L}_{t}^{\mu}f=\frac{1}{2}\mathrm{tr}(\sigma_{t}^{\mu}(\sigma_{t}^{\mu})^{\ast}\nabla^{2}f)+\nabla_{b_{t}^{\mu}}f,\ \ f\in C_{b}^{\infty}(\mathbb{R}^{d}),$
@@
Assume $\bf(A)$} and suppose that $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$ satisfies for some constant $\delta_{1}>0$,
(2.5) $\sigma_{t}(\gamma)\sigma_{t}(\gamma)^{\ast}\leq\delta_{1}I_{d\times d},\ \ \gamma\in\mathscr{P}_{2}(\mathbb{R}^{d}),t\geq 0.$
Then, $P_{s,t}^{\mu}$ satisfies the log-Sobolev inequality: for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$ and $f\in C_{b}^{\infty}(\mathbb{R}^{d})$, $f>0$,
(2.6) $P_{s,t}^{\mu}(f\log f)-P_{s,t}^{\mu}f\log P_{s,t}^{\mu}f\leq\frac{2\delta_{1}}{K_{1}}(\text{\rm{e}}^{K_{1}(t-s)}-1)P_{s,t}^{\mu}\big(|\nabla f^{\frac{1}{2}}|^{2}\big),\ \ 0\leq s\leq t.$
If moreover $\gamma\in\mathscr{P}_{2}(\mathbb{R}^{d})$ satisfies the log-Sobolev inequality: for some $L_{0}\geq 0$,
(2.7) $\gamma(f\log f)\leq\gamma(f)\log\gamma(f)+2L_{0}\gamma\big(|\nabla f^{\frac{1}{2}}|^{2}\big),\ \ f\in C_{b}^{\infty}(\mathbb{R}^{d}),f>0,$
then $P_{t}^{\ast}\gamma$ satisfies the log-Sobolev inequality:
(2.8) $\begin{split}(P_{t}^{\ast}\gamma)&(f\log f)-(P_{t}^{\ast}\gamma)(f)\log(P_{t}^{\ast}\gamma)(f)\\ &\leq 2\Big(L_{0}\text{\rm{e}}^{K_{1}t}+\frac{\delta_{1}}{K_{1}}(\text{\rm{e}}^{K_{1}t}-1)\Big)(P_{t}^{\ast}\gamma)\big(|\nabla f^{\frac{1}{2}}|^{2}\big),\ \ f\in C_{b}^{\infty}(\mathbb{R}^{d}),f>0,t\geq 0.\end{split}$
Remark 2.2.
In [35, Proposition 1.1], the authors derive the log-Sobolev inequality for the semigroup $P_{s,t}$ associated to
$\text{\rm{d}}X_{t}=b_{t}(X_{t})\text{\rm{d}}t+\text{\rm{d}}W_{t}$
under the assumption
$\langle b_{t}(x)-b_{t}(y),x-y\rangle\leq L|x-y|^{2}.$
Then, if the distribution $\mathscr{L}_{X_{0}}$ of the initial condition $X_{0}$ satisfies a log-Sobolev inequality, the distribution $\mathscr{L}_{X_{t}}$ at any $t>0$, also satisfies a log-Sobolev inequality (with a time-dependent constant).
Remark 2.3.
$\Gamma^{t,\mu}_{2}(f)\geq\rho\Gamma_{1}^{t,\mu}(f)=\rho|\nabla f|^{2}$
for some constant $\rho\in\mathbb{R}$ are equivalent, see e.g. [31, Proposition 3.12].
Remark 2.4.
One can also refer to [10, Theorem 4.1] for the log-Sobolev inequality of time-inhomogeneous semigroup and [40, Lemma 5.2] for the reflecting case under the Bakry-Émery curvature condition. Moreover, (2.10) is equivalent to [10, (10)] if we assume $\sigma\sigma^{\ast}\geq\delta_{2}$ for some $\delta_{2}>0$.
@@
For any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$, recall that $X_{s,t}^{\mu,x}$ solves (2.1) with initial value $x\in\mathbb{R}^{d}$ and $P_{s,t}^{\mu}$ is defined in (2.2). By ((A)) and $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$, we have
(2.9) $2\langle b_{t}(x,\mu_{t})-b_{t}(y,\mu_{t}),x-y\rangle\leq K_{1}|x-y|^{2},\ \ t\geq 0,x,y\in\mathbb{R}^{d}.$
It is easy to derive from (2.9) and $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$ that
$|X_{s,t}^{\mu,x}-X_{s,t}^{\mu,y}|\leq\text{\rm{e}}^{\frac{1}{2}K_{1}(t-s)}|x-y|,\ \ x,y\in\mathbb{R}^{d},0\leq s\leq t,$
which implies that for any $x\in\mathbb{R}^{d},0\leq s\leq t,f\in C_{b}^{\infty}(\mathbb{R}^{d})$,
@@
Next, we divide into three steps to complete the proof.
(i) We first assume (A), (2.5) and that for any $i\geq 1$, there exists a locally bounded function $\ell^{i}:[0,\infty)\rightarrow[0,\infty)$ such that for any $\mu_{\cdot}\in C([0,\infty),\mathscr{P}_{2}(\mathbb{R}^{d}))$,
(2.11) $\|\nabla^{i}b_{t}(\cdot,\mu_{t})\|_{\infty}\leq\ell^{i}_{t},\ \ t\geq 0,i\geq 1.$
(2.11) and $\sigma_{t}(x,\gamma)=\sigma_{t}(\gamma)$ imply that $P_{r,t}^{\mu}C_{b}^{\infty}(\mathbb{R}^{d})\subset C_{b}^{\infty}(\mathbb{R}^{d})$. By Itô’s formula, we conclude that
(2.12) $P_{r,t}^{\mu}f-P_{r,s}^{\mu}f=\int_{s}^{t}P_{r,\theta}^{\mu}\mathscr{L}_{\theta}^{\mu}f\text{\rm{d}}\theta,\ \ f\in C_{b}^{\infty}(\mathbb{R}^{d}),t\geq s\geq r,$
which together with the semigroup property implies that for $\varepsilon\geq 0$ and any $f\in C_{b}^{\infty}(\mathbb{R}^{d})$,
(2.13) $P_{r+\varepsilon,t}^{\mu}f-P_{r,t}^{\mu}f$ $=P_{r+\varepsilon,t}^{\mu}f-P_{r,r+\varepsilon}^{\mu}P_{r+\varepsilon,t}^{\mu}f=-\int_{r}^{r+\varepsilon}P_{r,\theta}^{\mu}\mathscr{L}_{\theta}^{\mu}P_{r+\varepsilon,t}^{\mu}f\text{\rm{d}}\theta.$
By (2.12), (2.13), (2.11) and $\sup_{t\in[0,T]}\|\sigma_{t}(\mu_{t})\|<\infty$ for any $T>0$, we may use the dominated convergence theorem to derive that for any $f\in C_{b}^{\infty}(\mathbb{R}^{d})$,
$-\int_{r}^{r+\varepsilon}P_{r,\theta}^{\mu}\mathscr{L}_{\theta}^{\mu}P_{r+\varepsilon,t}^{\mu}f\text{\rm{d}}\theta$
$=-\int_{r}^{r+\varepsilon}\mathscr{L}_{\theta}^{\mu}P_{\theta,t}^{\mu}f\text{\rm{d}}\theta+\int_{r}^{r+\varepsilon}(P_{r,r}^{\mu}\mathscr{L}_{\theta}^{\mu}P_{\theta,t}^{\mu}f-P_{r,\theta}^{\mu}\mathscr{L}_{\theta}^{\mu}P_{r+\varepsilon,t}^{\mu}f)\text{\rm{d}}\theta$
$=-\int_{r}^{r+\varepsilon}\mathscr{L}_{\theta}^{\mu}P_{\theta,t}^{\mu}f\text{\rm{d}}\theta+o(\varepsilon),\ \ f\in C_{b}^{\infty}(\mathbb{R}^{d}).$
This together with (2.13) implies the Kolmogorov backward equation, i.e.
(2.14) $\frac{\text{\rm{d}}P_{r,t}^{\mu}f}{\text{\rm{d}}r}=-\mathscr{L}_{r}^{\mu}P_{r,t}^{\mu}f,\ \ f\in C_{b}^{\infty}(\mathbb{R}^{d}),\ \ a.s.\ \ r\in[s,t].$